# Physics-informed, boundary-constrained Gaussian process regression for the reconstruction of fluid flow fields

### v1

@@ 5 Application to fluid flows around aerodynamic profiles
In this section, we perform GPR reconstructions of the velocity field of incompressible flows around aerodynamic profiles using the physics-informed framework presented above.
@@ 5.1 Profiles
    $y_{\text{c}}(x)$ =$=\begin{cases}\displaystyle\frac{\text{M}}{\text{P}^{2}}\left(2\text{P}x-x^{2}\right),&\text{for }0\leq x\leq\text{P}\,,\\ \displaystyle\frac{\text{M}}{(1-\text{P})^{2}}\left(1-2\text{P}+2\text{P}x-x^{2}\right),&\text{for }\text{P}<x\leq 1\,,\end{cases}$ (5.2)
@@ 5.3.1 Continuous profile boundary constraints
We first present the reconstruction of ${\boldsymbol{u}}$ by means of the physics-informed scalar kernel ${{G}_{0}}$ using only design points for the inlet and outlet boundaries ${{\boldsymbol{u}}_{\text{in}}}$ and ${{\boldsymbol{u}}_{\text{out}}}$, and for the boundary condition ${{\boldsymbol{u}}_{\text{w}}}={\boldsymbol{0}}$ on the tunnel walls. That is, we do not consider observation points in the interior of the domain in this reconstruction. The interpolated velocity field is presented in Figure 4. It can be observed that even though the information for flow reconstruction is insufficient at the interior of the computational domain compared to the ground truth, the slip boundary condition is already satisfied continuously along the profile boundary ${\Gamma_{\text{cyl}}}$. Moreover, in Figure 5 the computation of the total standard deviation of the velocity field (defined as the square root of the trace of the matrix-valued posterior covariance of the form (2.4)) for each approach is presented. We observe that constraining the base GP with the proposed BCGP-based approach around the profile boundary reduces uncertainty over this area, whereas the uncertainty is higher and uniform in the interior of ${\Omega}$ if only the unconstrained base kernel ${G}$ is used, since no physical information about the boundary condition is given.
@@
(b) Base scalar kernel ${G}$
Figure 5: Total standard deviation of the velocity field reconstructed by GPR with or without BCGP kernels. Design observations of the velocity are considered only at the outer boundary $\partial{\Omega}$ of the computational domain ${\Omega}$ ($N_{\text{b}}=116$, black circumferences $\circ$). (a) Estimate obtained by using the BCGP kernel ${{G}_{0}}$; (b) estimate obtained by using the base kernel ${G}$ instead of ${{G}_{0}}$, under the same observation design. The colormap represents the total standard deviation in m/s.
[…]
@@ 5.3.2 Discrete boundary constraints
To do this, we fix this budget in terms of the size of the Gram matrix used for interpolation, say ${N_{\text{Gram}}}=830$. Then, we give specific configurations of the design observations to each approach. For the BCGP-based method, we consider the same configuration of $N=415$ observations of the velocity as in Figure 6. In the case of the discrete approach, we consider $N=344$ observations of the velocity inside the domain ${\Omega}$ and its outer boundary, and we add ${N_{\Gamma}}=142$ observation points for the consideration of the slip boundary condition in Eq. (4.20). The velocity field estimates from physics-informed GPR with ${{G}_{0}}$ and with ${G}$ using discrete boundary constraints are presented in Figure 8. The relative error fields corresponding to each estimate, with respect to the ground truth from Figure 3, are also depicted. Furthermore, the configuration of the GPR design for each approach is summed up in Table 2, including the relative normal fit indicator $\epsilon_{n}$ of Eq. (4.15). It is observed that the BCGP-based method allows to satisfy the slip boundary condition at a better order of accuracy than that of the discrete method for the same observation budget. Moreover, this budget allows to include observations of the velocity at other parts of the domain instead of adding discrete points on the profile boundary.
@@
(b) Velocity field estimate using base kernel ${G}$.
@@
(d) Relative error field using base kernel ${G}$.
@@
GPR approach $N$ ${N_{\Gamma}}$ ${N_{\text{Gram}}}$ $\epsilon_{n}$
Scalar BCGP kernel ${{G}_{0}}$ 415 0 830 $2.780\cdot 10^{-5}$
Scalar base kernel ${G}$ 344 142 830 $1.611\cdot 10^{-2}$
Table 2: GPR design configurations and relative normal fit indicator $\epsilon_{n}$ of Eq. (4.15) for the GPR estimate of the velocity field around a cylinder profile using two approaches for the consideration of the slip boundary condition on the profile. The estimation using the scalar base kernel ${G}$ is obtained with discrete boundary constraints.
@@ 5.4 Flow around a NACA 0412 airfoil
This problem involves multi-scale dynamics of fluid motion. Therefore, estimates of the velocity field are constructed using an additive anisotropic Gaussian base kernel defined as:
${G}({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})=\sum_{m=1}^{M}\sigma_{m}^{2}\exp\left[-\frac{1}{2}\left(\frac{(x_{1}-x_{1}^{\prime})^{2}}{\ell_{m,1}^{2}}+\frac{(x_{2}-x_{2}^{\prime})^{2}}{\ell_{m,2}^{2}}\right)\right],\quad{\boldsymbol{x}},{\boldsymbol{x}}^{\prime}\in{\Omega}\,.$ (5.5)
This structure of the additive kernel is adapted to capture multiple scales of the flow dynamics [33]. In our context, we will consider only two scales ($M=2$) with the hyperparameters defined in Table 3. The computation of the associated BCGP kernel ${{G}_{0}}$ is done for each one of the choices for the measure ${\nu}$ from the Karhunen-Loève expansion over $\Gamma$, as in Corollaries 3.2 and 3.3.
Number of scales Large scale Small scale
$\sigma_{1}=10$ $\sigma_{2}=0.005^{4}$
$M=2$ $\ell_{1,1}=0.6$ $\ell_{2,1}=0.01$
$\ell_{1,2}=0.08$ $\ell_{2,2}=0.005$
Table 3: Hyperparameters of the two-scale additive base kernel used for velocity GPR estimates of the fluid flow around the NACA 0412 airfoil.
The GPR observation design for each of the two approaches is done following a budget fixed by the Gram matrix size as ${N_{\text{Gram}}}=780$. In the BCGP-based case, we considered $N=390$ observations of velocity inside the domain, while in the discrete approach, we fixed $N=370$ positions for velocity information and ${N_{\Gamma}}=40$ discrete points along the boundary of the airfoil leading edge. These configurations along with the relative normal fit indicator from Eq. (4.15) are displayed in Table 4 for each approach. The nugget regularization parameter is fixed as $\eta=1\cdot 10^{-7}$. In Figure 11 a convergence study for the satisfaction of the profile boundary condition in terms of indicators from Eq. (4.14) and Eq. (4.15) with respect to the spectral accuracy of the kernel truncation is presented. Both choices of ${\nu}$ are taken into account for comparison. The numerical parameters for obtaining these estimates are $I=2000$ and $J+1=55$ (see Algorithms 1 and 2).
It is observed that under this specific configuration of kernel and profile geometry, the relative accuracy for the fulfillment of the slip boundary condition is of the same order among the two methods. However, in the BCGP-based case, there is no GPR design which involves discrete observation of this condition in the posterior estimate.
@@
GPR approach $N$ ${N_{\Gamma}}$ ${N_{\text{Gram}}}$ $\epsilon_{n}$
Base kernel ${G}$ 370 40 780 $5.61\cdot 10^{-3}$
BCGP kernel ${{G}_{0}}$ (pushforward ${\nu}$) 390 0 780 $4.62\cdot 10^{-4}$
BCGP kernel ${{G}_{0}}$ (surface ${\nu}$) 390 0 780 $5.14\cdot 10^{-4}$
Table 4: GPR design configurations and relative normal fit indicator $\epsilon_{n}$ from Eq. (4.15) for the GPR estimate of the velocity field around the leading edge of a NACA 0412 airfoil using multiple approaches for the inclusion of the slip boundary condition on the profile. The first approach considers the base kernel ${G}$ and ${N_{\Gamma}}=40$ discrete points for the observation of the slip boundary condition around the profile boundary. The other two approaches consider the associated BCGP kernel ${{G}_{0}}$ instead. Each one of them corresponds to a different choice of the measure ${\nu}$ from the Karhunen-Loève expansion, as described in Corollaries 3.2 and 3.3.
@@ 6 Conclusions and perspectives
In this work, we have developed an approach based on Gaussian process (GP) regression for simulating incompressible fluid flow fields around aerodynamic profiles by defining physics-informed GP priors. This reconstruction method is meshfree, geometrically flexible, and is intended to provide estimates from scarce data scenarios by informing the prior distributions of the flow fields with their underlying physics. These priors thereby satisfy a divergence-free condition for the velocity field by modelling it as a linear differential operator of a stream function which is itself described by a scalar GP. This process is in turn constrained on the boundary of an aerodynamic profile inside the domain, such that the corresponding velocity estimate satisfies a slip boundary condition on that profile. Hence the divergence-free condition is enforced in a continuous way everywhere inside the computational domain, and the slip condition is enforced all along the profile boundary alike.

### v2

@@ 5 Application to fluid flows around aerodynamic profiles
In this section, we perform GPR reconstructions of the velocity field of incompressible flows around aerodynamic profiles using the physics-informed framework presented above. The numerical implementation for the presented results on the cylinder profile and NACA 0412 airfoil is available at https://github.com/adrirps/Physics-Informed-BCGPs-for-Flow-Fields-Reconstruction. This Python implementation uses symbolic computation, through the package SymPy [32], for getting high precision in kernel computations, including derivatives.
@@ 5.1 Profiles
[…]
@@ 5.3.1 Continuous profile boundary constraints
We first present the reconstruction of $\boldsymbol{u}$} by means of the physics-informed scalar kernel ${G}_{0}$} using only design points for the inlet and outlet boundaries ${\boldsymbol{u}}_{\text{in}}$} and ${\boldsymbol{u}}_{\text{out}}$}, and for the boundary condition ${{\boldsymbol{u}}_{\text{w}}}={\boldsymbol{0}}$ on the tunnel walls. That is, we do not consider observation points in the interior of the domain in this reconstruction. The interpolated velocity field is presented in Figure 4. It can be observed that even though the information for flow reconstruction is insufficient at the interior of the computational domain compared to the ground truth, the slip boundary condition is already satisfied continuously along the profile boundary $\Gamma_{\text{cyl}}$}. Moreover, in Figure 5 the computation of the total standard deviation of the velocity field (defined as the square root of the trace of the matrix-valued posterior covariance of the form (2.4)) for each approach is presented. We observe that constraining the base GP with the proposed BCGP-based approach around the profile boundary reduces uncertainty over this area, whereas the uncertainty is higher and uniform in the interior of $\Omega$} if only the unconstrained base kernel G{G} is used, since no physical information about the boundary condition is given.
@@
(b) Base scalar kernel G{G}
Figure 5: Total standard deviation of the velocity field reconstructed by GPR with or without BCGP kernels. Design observations of the velocity are considered only at the outer boundary $\partial{\Omega}$ of the computational domain $\Omega$} ($N_{\text{b}}=116$, black circumferences $\circ$). (a) Estimate obtained by using the BCGP kernel ${G}_{0}$}; (b) estimate obtained by using the base kernel G{G} instead of ${G}_{0}$}, under the same observation design. The colormap represents the total standard deviation in m/s.
Next, we present the GPR reconstruction of the velocity field using our physics-informed framework by adding observation points inside the computational domain. That is, we consider the reconstruction formula (4.9) with ${\boldsymbol{X}}={{\boldsymbol{X}}_{\text{b}}}\cup{{\boldsymbol{X}}_{\text{d}}}$ and ${\boldsymbol{V}}(t)=\smash{({\boldsymbol{V}}_{\text{b}}(t)^{\mathsf{T}},{\boldsymbol{V}}_{\text{d}}(t)^{\mathsf{T}})}^{\mathsf{T}}$ as in Sect. 4.3. A total of $N=415$ observation points of the velocity was considered in the computational domain $\Omega$}, of which $N_{\text{b}}=88$ and ${N_{{\Omega}}}=327$ observations belong to $\partial{\Omega}$ and $\Omega$}, respectively. As before, no design observation on the profile boundary $\Gamma_{\text{cyl}}$} is included. In order to compute the spectral expansion for the definition of $\smash{{{G}_{0}^{(I,J)}}}$ from Eq. (3.23) and (4.19), a uniform regular discretization of ${\mathcal{P}}=[0,2\pi]$ in $I=400$ intervals and the largest $J+1=19$ eigenvalues of the expansion were used (see below). The interpolation of the velocity field is presented in Figure 6.
@@ 5.3.2 Discrete boundary constraints
To do this, we fix this budget in terms of the size of the Gram matrix used for interpolation, say ${N_{\text{Gram}}}=830$. Then, we give specific configurations of the design observations to each approach. For the BCGP-based method, we consider the same configuration of $N=415$ observations of the velocity as in Figure 6. In the case of the discrete approach, we consider $N=344$ observations of the velocity inside the domain $\Omega$} and its outer boundary, and we add ${N_{\Gamma}}=142$ observation points for the consideration of the slip boundary condition in Eq. (4.20). The velocity field estimates from physics-informed GPR with ${G}_{0}$} and with G{G} using discrete boundary constraints are presented in Figure 8. The relative error fields corresponding to each estimate, with respect to the ground truth from Figure 3, are also depicted. Furthermore, the configuration of the GPR design for each approach is summed up in Table 2, including the relative normal fit indicator $\epsilon_{n}$ of Eq. (4.15). It is observed that the BCGP-based method allows to satisfy the slip boundary condition at a better order of accuracy than that of the discrete method for the same observation budget. Moreover, this budget allows to include observations of the velocity at other parts of the domain instead of adding discrete points on the profile boundary.
@@
(b) Velocity field estimate using base kernel G{G}.
@@
(d) Relative error field using base kernel G{G}.
@@
GPR approach NN $N_{\Gamma}$} $N_{\text{Gram}}$} $\epsilon_{n}$
Scalar BCGP kernel ${G}_{0}$} 415 0 830 $1.823\cdot 10^{-5}$
Scalar base kernel G{G} 344 142 830 $1.611\cdot 10^{-2}$
Table 2: GPR design configurations and relative normal fit indicator $\epsilon_{n}$ of Eq. (4.15) for the GPR estimate of the velocity field around a cylinder profile using two approaches for the consideration of the slip boundary condition on the profile. The estimation using the scalar base kernel G{G} is obtained with discrete boundary constraints.
@@ 5.4 Flow around a NACA 0412 airfoil
Since this problem involves multi-scale dynamics of fluid motion, we define an additive anisotropic Gaussian kernel of the form:
${G}({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})=\sum_{m=1}^{M}\sigma_{m}^{2}\exp\left[-\frac{1}{2}\left(\frac{(x_{1}-x_{1}^{\prime})^{2}+\alpha_{m}^{2}(x_{2}-x_{2}^{\prime})^{2}}{\ell_{m}^{2}}\right)\right],\quad{\boldsymbol{x}},{\boldsymbol{x}}^{\prime}\in{\Omega}\,.$ (5.5)
In our context, we consider this structure with two scales ($M=2$) as the base kernel for boundary-constraining. In particular, we set $\ell_{1}=0.64$ and $\alpha_{1}=8$ in the larger scale, while we fix $\ell_{2}=\ell_{1}/2^{6}$ and $\alpha_{2}=2$ to capture the motion in the smaller scale. We then set $\sigma_{2}=\sigma_{1}/2^{12}$ with $\sigma_{1}=0.4$ to adapt the kernel to the fluid energy decay, as explained in [34]. The computation of the associated BCGP kernel ${G}_{0}$} is done for each one of the choices of the measure $\nu$} from the Karhunen-Loève expansion over $\Gamma$, as in Corollaries 3.2 and 3.3.
The GPR observation design for each of the two approaches is done following a budget fixed by the Gram matrix size as ${N_{\text{Gram}}}=780$. In the BCGP-based case, we considered $N=390$ observations of velocity inside the domain, while in the discrete approach, we fixed $N=370$ positions for velocity information and ${N_{\Gamma}}=40$ discrete points along the boundary of the airfoil leading edge. These configurations along with the relative normal fit indicator from Eq. (4.15) are displayed in Table 3 for each approach. The nugget regularization parameter is fixed as $\eta=1\cdot 10^{-10}$. In Figure 11 a convergence study for the satisfaction of the profile boundary condition in terms of indicators from Eq. (4.14) and Eq. (4.15) with respect to the spectral accuracy of the kernel truncation is presented. Both choices of $\nu$} are taken into account for comparison. The numerical parameters for obtaining these estimates are $I=2000$ and $J+1=78$ (see Algorithms 1 and 2).
It is observed that under this specific configuration of kernel and profile geometry, the relative accuracy for the fulfillment of the slip boundary condition is better using BCGPs. Moreover, in this case, there is no GPR design which involves discrete observation of this condition in the posterior estimate.
@@
GPR approach NN $N_{\Gamma}$} $N_{\text{Gram}}$} $\epsilon_{n}$
Base kernel G{G} 370 40 780 $5.76\cdot 10^{-3}$
BCGP kernel ${G}_{0}$} (pushforward $\nu$}) 390 0 780 $6.76\cdot 10^{-5}$
BCGP kernel ${G}_{0}$} (surface $\nu$}) 390 0 780 $8.02\cdot 10^{-5}$
Table 3: GPR design configurations and relative normal fit indicator $\epsilon_{n}$ from Eq. (4.15) for the GPR estimate of the velocity field around the leading edge of a NACA 0412 airfoil using multiple approaches for the inclusion of the slip boundary condition on the profile. The first approach considers the base kernel G{G} and ${N_{\Gamma}}=40$ discrete points for the observation of the slip boundary condition around the profile boundary. The other two approaches consider the associated BCGP kernel ${G}_{0}$} instead. Each one of them corresponds to a different choice of the measure $\nu$} from the Karhunen-Loève expansion, as described in Corollaries 3.2 and 3.3.
@@ 6 Conclusions and perspectives
In this work, we have developed an approach based on Gaussian process (GP) regression for simulating incompressible fluid flow fields around aerodynamic profiles by defining physics-informed GP priors. This reconstruction method is mesh-free, geometrically flexible, and is intended to provide estimates from scarce data scenarios by informing the prior distributions of the flow fields with their underlying physics. These priors thereby satisfy a divergence-free condition for the velocity field by modelling it as a linear differential operator of a stream function which is itself described by a scalar GP. This process is in turn constrained on the boundary of an aerodynamic profile inside the domain, such that the corresponding velocity estimate satisfies a slip boundary condition on that profile. Hence the divergence-free condition is enforced in a continuous way everywhere inside the computational domain, and the slip condition is enforced all along the profile boundary alike.